OpenStudy (anonymous):

Simplify: (x^2 - 9)/(x^3 - 3x^2)

OpenStudy (jim766):

factor the numerator ,the top factor the denominator, the bottom any idea how to start?

OpenStudy (anonymous):

I know how to do it but could you answer it since this one just confuses me

OpenStudy (jim766):

x^2-9 is a difference of 2 squares do you know how that factors?

OpenStudy (anonymous):

\[\frac{x^2-9}{x^3 -3x^2} \\ \frac{ ( x-3)(x+3)}{x^2( x -3)}\\ \frac{ ( x+3)}{x^2}\\\]